OpenStudy (ledah):

Solve the absolute value equation |8n + 4| = 44. One solution is n = 5. What is the other solution? n = -5 n = -6 n = 6 none of the above

OpenStudy (anonymous):

8n+4=44 or 8n+4=-44... first one gives 5 and second one gives -6

OpenStudy (theviper):

n=5

OpenStudy (theviper):

none of the above

OpenStudy (anonymous):

even yours?

OpenStudy (jackellyn):

8n +4 =44 8n = 40 n = 5 8n +4 = -44 8n = -48 n =-6

OpenStudy (anonymous):

|8n + 4| = 44 8n + 4 = 44 8n + 4 = -44 -4 -4 -4 -4 8n = 40 8n = -48 /8 /8 /8 /8 n = 5 n = -6 So, it would "n = -6" =p

OpenStudy (anonymous):

1) you have to remember that equations like this will always have two values 2) the first case will be 8n+4=44 3) then you subtract 4 and divide by eight to get n=5 4) the second case would be 8n+4=-44 5) you then subtract from both sides and divide to get n=-6 hope that helps :D

OpenStudy (anonymous):

none of the above

OpenStudy (ledah):

are you guys even trying or just giving random answers....?

OpenStudy (theviper):

best answer by everyone

OpenStudy (anonymous):

ledah i am actually trying to help you out on this, not giving you random answers